A parallel plate capacitor of plate area A square meters is connected to a 10-volt battery. The plates are moved apart with constant speed v0, while still connected to the 10-volt battery. Find the instantaneous rate of change of potential energy stored in a capacitor when the plates are x meters apart.
Answer #1

Energy = 1/2 CV^2

C= A *permittivity of free space/distance between capacitor plates

instantaneous rate of change of potential energy =dE/dt = d(1/2 V^2 x C)/dt

= 0.5 x 10^2 x dC/dt= 50x A x permittivity of free space/x dt=50 x permittivity of free space /v0